Salsa Inc Income Statement For the month ended August 31,20X1 Service revenue Expenses Salaries expense Rent expense Interest expense Utilities expense Other expense $ 50,000 10,000 5,000 4,000 8,000 Total expenses Net Income $_200,000 Salsa Inc. Statement of Stockholders Equity For the month ended August 31,20X1 Common Stock $ 20,000 8,000 Retained Earnings $ 19,000 Total Stockholders Equity Beginning balances - Aug 1 Add: New stock issuances Add: Net income Less: Dividends $ 39,000 8,000 200,000 200,000 Ending balances - Aug 31 S 28,000 $213.000 241000 Salsa Inc Balance Sheet at August 31,20X1A. From the balance sheet, what is the cash account balance at Aug 31? Answer: S B. From the Balance Sheet, what is the Retained Earnings balance? (Hint: this answer should also match the Retained Earnings at Aug 31 from the Statement of Stockholders Equity Answer: S C. From the Statement of Stockholders Equity, how much in dividends were paid for the month of August? Answer: D. From the Income Statement, how much service revenue was earned for the month of August? Answer: $
Answer #1

Answers

  • Answer ‘A’

Cash balance = $ 46,000

Working:
Total Liabilities & Stockholder’s Equity = Total Assets
$ 274000 = Total Assets
$ 274000 = Cash + Other Assets
274000 = Cash + 35000 + 13000 + 60000 + 120000
274000 – 35000 – 13000 – 60000 – 120000 = CASH
Cash = $ 46,000

  • Answer ‘B’

Retained Earnings balance on Aug 31 = $ 213,000
For answer, have a look at Statement of Stockholder’s Equity.
In there, check out the TOTAL (or Ending balances on Aug 31) of Retained earnings for answer.

  • Answer ‘C’

Dividend paid = $ 6,000
For this answer, go back to Statement of Stockholder’s Equity.
The Dividend amount is missing, but total $ 213000 is given.
Hence, $ 19000 + $ 200000 net income – Dividends = $ 213000 ending balance
219000 – Dividends = 213000
219000 – 213000 = Dividends
Dividend = $ 6000

  • Answer ‘D’

Total Service Revenue = $ 277,000

Revenue earned = Net Income + Expenses
= 200000 + 50000 + 10000 + 5000 + 4000 + 8000
= $ 277,000